OpenStudy (fanduekisses):

How do I find the zeros of 3x^3-x^2-2x?

OpenStudy (fanduekisses):

\[3x ^{3}-x ^{2}-2x\]

OpenStudy (fanduekisses):

x(3x^2-x-2) ? I'm stuck :(

OpenStudy (campbell_st):

ok... so you now need to factor the quadratic... so looking at \[3x^2 - x - 2\] multiply 3 and -2 then find the factors that add to -1 the larger factor is negative.... then its (3x + factor 1)(3x + factor 2) -------------------------- 3 remove a common factor from a binomial, and you'll find it cancels with the denominator... whats left is the factored form... hope it helps

OpenStudy (mathmale):

Review: "zeros" are values of x at which the given function is zero. Fanduekisses has done a great job of typing out this expression in proper math symbols. How would you factor \[3x^2 -x -2?\] The first thing I'd do would be to multiply the first and last coefficients of this polynomial together: 3*2=-6. Then I'd list all integer factors of the result (6), to obtain -6*1 -3*2 -2*3 and so on. Next, I'd look for the pair of factors, if any, whose sum is -1, the coeff. of the x term.

OpenStudy (mathmale):

Which pair would that be?

OpenStudy (fanduekisses):

-3 and 2

OpenStudy (mathmale):

Campbell and I have presented essentially the same approach. According to rules of mathematical logic, does that mean that he and I are both right? :)

OpenStudy (mathmale):

If you choose -3 and 2, which i would choose myself, then I'd "factor by grouping:" 3x^2 -x - 2 becomes 3x^2 -3x + 2x - 2, which follows your choice of -3 and 2 as factors of -6 Factor by grouping: 3x(x-1) + 2(x-1) x-1 is the factor common to both sides, so factor that out and determine what remains; what remains is the 2nd factor. Can you finish this?

OpenStudy (fanduekisses):

ohh yes :) so zeros are 1 and 1?

OpenStudy (mathmale):

Campbell: What would you do here to determine the zeros?

OpenStudy (campbell_st):

lol... oh well (3x -3)(3x + 2) ------------- 3 remove a common factor from the 1st binomial 3(x -1)(3x +2) ------------- 3 cancel the common factor leaves (x -1)(3x +2) so the factored form is x(x-1)(3x+2) the reason why its confusing is teachers make students split the middle term and group in pairs... which can be extremely difficult for students...

OpenStudy (fanduekisses):

oooohhhhh I see, yeah at first I found it confusing lol yeah some teachers make us do it differently which is actually more confusing :) Thanks guys :)

OpenStudy (campbell_st):

to find the zeros set each factor to zero x = 0 x - 1 = 0 3x + 2 = 0 and solve each for x

OpenStudy (mathmale):

Hope that you, Fondue, will check your answers. Substitute each supposed zero back into the original expression. If the expression is then zero, great; you have that zero correct. Check the other supposed zero also.

OpenStudy (fanduekisses):

Thanks so much guys <3

OpenStudy (mathmale):

Very happy to be of help. Thanks again, Campbell.